an electron moving perpendicular to a uniform magnetic field of 0.22 T moves in a circle with a speed of 1.5×10^7 m/s. what is the radius of the circle? (e= 1.60×10^-10^-19 C, M electron =9.11×10^-31 kg)
Answer #1

here,

magnetic feild , B = 0.22 T

speed , v = 1.5 * 10^7 m/s

the radius of circle , R = m * v /( q * B)

R = 9.1 * 10^-31 * 1.5 * 10^7 /( 1.6 * 10^-19 * 0.22) m

R = 3.89 * 10^-2 m

the radius of circle is 3.89 * 10^-2 m